OpenStudy (marsmcnugget):

What is the x-intercept of the line with the equation -2x+1/2y=18 ? ( , 0)

OpenStudy (marsmcnugget):

@Rushwr

OpenStudy (marsmcnugget):

@ytrewqmiswi

OpenStudy (marsmcnugget):

@raiders88

OpenStudy (onepieceftw):

The x-intercept is the x value when y=0 -2x+1/2(0)=18 Solve that ^

OpenStudy (marsmcnugget):

is the answer 3?

OpenStudy (marsmcnugget):

@OnePieceFTW

OpenStudy (onepieceftw):

\[-2x + \frac{ 1 }{ 2 }(0) = 18\]

OpenStudy (onepieceftw):

What is 1/2 times 0?

OpenStudy (marsmcnugget):

0

OpenStudy (onepieceftw):

So no we have -2x=18 What do we do next?

OpenStudy (onepieceftw):

now*

OpenStudy (marsmcnugget):

the answer is -9 right?

OpenStudy (marsmcnugget):

@OnePieceFTW

OpenStudy (onepieceftw):

Yes, that is correct.

OpenStudy (onepieceftw):

Good work :)

OpenStudy (marsmcnugget):

thanks :D